# Editable-DeepSC: Reliable Cross-Modal Semantic Communications for Facial Editing

### v3

@@ 5.1 Experiment Setup
(a) LPIPS Results
(b) FID Results
(c) KID Results
(d) CLIPsim Results
@@
Simulation Datasets. As for the textual instruction $S_{T}$, we use the editing requests from the CelebA-Dialog dataset [43] for all the experiments. As for the original image $S_{I}$: (1) In the main experiments, we consider relatively simple scenarios and choose from the CelebA dataset [46], whose images are cropped to the resolution of $128\times 128$ and have similar distribution with the training set of StyleGAN [25]; (2) In the more realistic high-resolution experiments, we choose from the CelebA-HQ dataset [47] with the resolution of $1024\times 1024$; (3) In the more realistic out-of-distribution (OOD) experiments, we choose from the MetFaces dataset [48] with the resolution of $128\times 128$, whose facial images are from artworks and have an inherent distribution bias with the natural facial images of StyleGAN training set. As for the fine-tuning images, we use images from the CelebA dataset [46] for the resolution of $128\times 128$, and use images from the CelebA-HQ dataset [47] for the resolution of $1024\times 1024$. We strictly ensure that the fine-tuning images are few in number ($n_{ft}<20$) and have no intersections with the images used in the test time.
Implementation Details. To simulate the varying communication circumstances, we set the SNR levels of 6-6 dB, 0 dB, $6$ dB, $12$ dB, $18$ dB for all the methods. For Editable-DeepSC, $\lambda_{inv}$, $\eta_{inv}$, $r_{inv}$, $\eta_{ft}$, $r_{ft}$, and $n_{ft}$ are set as $1.0$, $0.10$, $110$, $1\times 10^{-4}$, $6$, and $10$, respectively. For DeepSC [5], the learning rate is set as $1\times 10^{-3}$ and the batch size is set as $4$. For JPEG [49] and LDPC [50], the target bpp under the resolution of $128\times 128$ is set as $0.7$, the target bpp under the resolution of $1024\times 1024$ is set as $0.2$, and the QAM order is set as $4$. For Talk2Edit [22], we adopt its official code implementation. To ensure a fair comparison, the learning-based transmission approach, DeepSC [5], is first fully trained under each considered SNR level before testing. All the experiments are conducted on NVIDIA RTX A6000 GPUs.
5.2 Main Results
@@
(a) LPIPS Results
(b) FID Results
(c) KID Results
(d) CLIPsim Results
@@
We compare Editable-DeepSC with the baselines on the CelebA dataset. From Figure 4, we find that Editable-DeepSC realizes the best results in terms of LPIPS, FID, KID, and CLIPsim with significant improvements. For instance, when the SNR level is 6-6 dB, the LPIPS, FID, KID, and CLIPsim scores of Editable-DeepSC are improved by $47.65\%$, $67.82\%$, $101.05\%$, and $24.76\%$ than the second best method, respectively. These results indicate that Editable-DeepSC achieves excellent editing effects with high fidelity and quality, outperforming the baselines that separately handle communications and editings. We also observe that the performance of JPEG+LDPC+Talk2Edit and Talk2Edit+JPEG+LDPC will decrease rapidly when the SNR level is below $6$ dB, which is consistent with the fact that the traditional communications suffer greatly from the cliff effect [26].
Figure 5 illustrates the qualitative comparison of different methods at the noise level of $6$ dB. We notice that Editable-DeepSC indeed achieves the best visualization results. Although the other methods also manage to edit the images, their effects are not as vivid and natural as those achieved with Editable-DeepSC. This is because Editable-DeepSC reduces the data processing procedures and preserves more semantic mutual information, which eventually leads to more satisfying editing performance. Table 1 shows the compression effectiveness of different methods from the perspective of CBR defined in (5). We find that Editable-DeepSC only utilizes around $12.5\%$ or $26.7\%$ of the baselines’ CBR, yet it still achieves extraordinary editing effects and outperforms all the baselines. Editable-DeepSC not only performs high-quality editings, but also considerably saves the transmission bandwidth.
@@
(a) LPIPS Results
(b) FID Results
(c) KID Results
(d) CLIPsim Results
@@
5.3 High-Resolution Scenario Results
We then consider more realistic settings where the image resolution is expanded to $1024\times 1024$ on the CelebA-HQ dataset. Figure 6 presents the LPIPS, FID, KID, and CLIPsim results while Table 2 presents the CBR results. The amplification of image pixels objectively increases the complexity of editing and transmission tasks. However, we conclude that Editable-DeepSC again achieves the best LPIPS scores under almost all the tested cases. As for the FID, KID, and CLIPsim scores, although Editable-DeepSC can’t obtain the best performance in some cases, it only uses around $0.2\%$ or $1.8\%$ of the baselines’ CBR with much lower bandwidth consumption as shown in Table 2. In general, Editable-DeepSC still exhibits great robustness and efficiency under high-resolution scenarios.
@@ 5.4 Out-Of-Distribution (OOD) Scenario Results
The visualization results in Figure 9 further demonstrate that Editable-DeepSC can precisely edit the OOD images without influencing other unrelated regions. In contrast, all the other methods struggle to realize the semantic editing while maintaining the fidelity, and encounter image quality degradation or even collapse under the more rigorous OOD scenarios. Our combination of editings and communications has indeed come into effect for better downstream tasks performance.
5.5 Further Analysis
The influences of pre-trained GAN inversion. First, we explore the impacts of the proposed GAN inversion based semantic coding by changing the number of total inversion iterations $r_{inv}$. From Table 3, we notice that when $r_{inv}$ is small, increasing its value indeed improves the editing performance as the GAN inversion is more fully optimized, which validates the contribution of this component. However, when $r_{inv}$ is large, further increasing its value cannot improve the editing performance and will also introduce more GPU energy costs. Thus, our adoption of $r_{inv}=110$ is reasonable, for it strikes a better balance between editing effects and GPU energy costs.
The influences of iterative attributes matching. Next, we study the impacts of the iterative attributes matching mechanism by removing it from Editable-DeepSC. As shown in Figure 11, when the iterative attributes matching mechanism is excluded, Editable-DeepSC will be unable to precisely perform the anticipated editings according to the given texts in a fine-grained way. This again provides evidence for the contribution of our iterative attributes matching mechanism.
The influences of model fine-tuning. Similarly, we also investigate the impacts of the model fine-tuning mechanism in Table 3, Table 5, and Figure 11. We discover that under the extreme channel condition of 6-6 dB SNR, the fine-tuning mechanism plays a crucial role in preserving the quality and fidelity of edited images, while removing it will conversely generate unnatural or even collapsed regions. These results also verify the contribution of our fine-tuning mechanism. Moreover, our adoption of $r_{ft}=6$ yields a satisfactory trade-off between fine-tuning effects and GPU energy consumption.
Table 3: Quantitative results under different $r_{inv}$ and $r_{ft}$ on the CelebA dataset (resolution $128\times 128$) at the SNR of 6-6 dB.
Table 5: GPU energy costs of different methods averaged over each image under different resolutions.
@@ 6 Conclusion
In this paper, we propose Editable-DeepSC, a novel cross-modal semantic communication approach that tackles the communication challenges under dynamic facial editing scenarios in a conversational and interactive way. Specifically, we first theoretically analyze different transmission schemes that separately handle communications and editings. We find that they actually increase the data processing procedures and are more likely to lose semantic mutual information. In light of this, we emphasize the necessity to integrate editings into the communication chain, jointly optimize these two aspects, and preserve more semantic mutual information. To compress the high-dimensional data, we leverage inversion methods based on pre-trained StyleGAN priors for semantic coding. To tackle the dynamic channel noise conditions, we propose SNR-aware channel coding based on model fine-tuning, which introduces two lightweight trainable adapters to capture the distribution of new noise conditions. Extensive experiments demonstrate that compared to existing methods, Editable-DeepSC exhibits superior editing effects while significantly saving the transmission bandwidth under multiple settings.

### v4

@@ 5 Experiments
In this section, we conduct comprehensive experiments to evaluate the performance of the proposed Editable-DeepSC framework. First, the experimental setup (including datasets, baselines, and evaluation metrics) is introduced. The main results on the standard-resolution CelebA dataset [30] are then presented, followed by evaluations under more challenging high-resolution and out-of-distribution (OOD) scenarios. Finally, we provide detailed ablation studies and further analyses to verify the effectiveness of the key components and to gain deeper insights behind the method design.
@@ 5.1 Experiment Setup
Simulation Datasets. As for the textual instruction $S_{T}$, we use the editing requests from the CelebA-Dialog dataset [23] for all the experiments. As for the original image $S_{I}$: (1) In the main experiments, we consider relatively simple scenarios and choose from the CelebA dataset [30], whose images are cropped to the resolution of $128\times 128$ and have similar distribution with the training set of StyleGAN [27]; (2) In the more realistic high-resolution experiments, we choose from the CelebA-HQ dataset [25] with the resolution of $1024\times 1024$; (3) In the more realistic out-of-distribution (OOD) experiments, we choose from the MetFaces dataset [26] with the resolution of $128\times 128$, whose facial images are from artworks and have an inherent distribution bias with the natural facial images of StyleGAN training set. As for the fine-tuning images, we use images from the CelebA dataset [30] for the resolution of $128\times 128$, and use images from the CelebA-HQ dataset [25] for the resolution of $1024\times 1024$. We strictly ensure that the fine-tuning images are few in number ($n_{ft}<20$) and have no intersections with the tested images.
Implementation Details. To simulate the varying communication circumstances, we set the SNR levels of 6-6 dB, 0 dB, $6$ dB, $12$ dB, $18$ dB for all the methods. For Editable-DeepSC, $\lambda_{inv}$, $\eta_{inv}$, $r_{inv}$, $\eta_{ft}$, $r_{ft}$, and $n_{ft}$ are set as $1.0$, $0.10$, $110$, $1\times 10^{-4}$, $6$, and $10$, respectively. For DeepSC [58], the learning rate is set as $1\times 10^{-3}$ and the batch size is set as $4$. For JPEG [43] and LDPC [14], the target bpp under the resolution of $128\times 128$ is set as $0.7$, the target bpp under the resolution of $1024\times 1024$ is set as $0.2$, and the Quadrature Amplitude Modulation (QAM) order is set as $4$. For Talk2Edit [24], we adopt its official code implementation. To ensure a fair comparison, the learning-based transmission approach, DeepSC [58], is first fully trained under each considered SNR level before testing. All the experiments are conducted on NVIDIA RTX A6000 GPUs.
5.2 Main Results
We compare Editable-DeepSC with the baselines on the CelebA dataset. From Figure 4, we find that Editable-DeepSC realizes the best results in terms of LPIPS, FID, KID, and CLIPsim with significant gains. For instance, when the SNR level is 6-6 dB, the LPIPS, FID, KID, and CLIPsim scores of Editable-DeepSC are improved by $47.65\%$, $67.82\%$, $101.05\%$, and $24.76\%$ than the second best method, respectively. These results indicate that Editable-DeepSC achieves excellent editing effects with high fidelity and quality, outperforming the baselines that separately handle communications and editings. We also observe that the performance of JPEG+LDPC+Talk2Edit and Talk2Edit+JPEG+LDPC will decrease rapidly when the SNR level is below $6$ dB, which is consistent with the fact that the traditional communications suffer greatly from the cliff effect [5].
Figure 5 illustrates the qualitative comparison of different methods at the noise level of $6$ dB. We notice that Editable-DeepSC indeed achieves the best visualization results. Although the other methods also manage to edit the images, their effects are not as vivid and natural as those achieved with Editable-DeepSC. This is because Editable-DeepSC reduces the data processing procedures and preserves more semantic mutual information, which eventually leads to more satisfying editing results. Table 2 shows the compression effectiveness of different methods in terms of CBR defined in (5). We find that Editable-DeepSC only utilizes around $12.5\%$ or $26.7\%$ of the baselines’ CBR, yet it still achieves extraordinary editing effects and outperforms all the baselines. Editable-DeepSC not only performs high-quality editings, but also considerably saves the transmission bandwidth.
5.3 High-Resolution Scenario Results
We then consider more realistic settings where the image resolution is expanded to $1024\times 1024$ on the CelebA-HQ dataset. Figure 6 presents the LPIPS, FID, KID, and CLIPsim results while Table 3 presents the CBR results. The amplification of image pixels objectively increases the complexity of editing and transmission tasks. However, we find that Editable-DeepSC again achieves the best LPIPS scores under almost all the tested cases. As for the FID, KID, and CLIPsim scores, although Editable-DeepSC can’t obtain the best results in some cases, it only uses around $0.2\%$ or $1.8\%$ of the baselines’ CBR with much lower bandwidth costs as shown in Table 3. In general, Editable-DeepSC still exhibits great robustness and efficiency under high-resolution scenarios.
@@ 5.4 Out-Of-Distribution (OOD) Scenario Results
The visualization results in Figure 9 further demonstrate that Editable-DeepSC can precisely edit the OOD images without influencing other unrelated regions. In contrast, all the other methods struggle to realize the semantic editing while maintaining the fidelity, and encounter image quality degradation or even collapse under the more rigorous OOD scenarios. Our combination of editings and communications has indeed come into effect for better tasks performance.
@@
5.5 Further Analysis
Table 4: Quantitative results under different values of $r_{inv}$ and $r_{ft}$ on the CelebA dataset (resolution $128\times 128$) at the SNR of 6-6 dB.
The influences of pre-trained GAN inversion. First, we explore the impacts of the proposed GAN inversion based semantic coding by changing the value of $r_{inv}$. From Table 4, we notice that when $r_{inv}$ is small, increasing its value indeed improves the editing performance as the GAN inversion is more fully optimized, which validates the contribution of this component. However, when $r_{inv}$ is large, further increasing its value cannot improve the editing performance and will also introduce more GPU energy costs. Thus, our adoption of $r_{inv}=110$ is reasonable, for it strikes a better balance between editing effects and GPU energy costs.
The influences of iterative attributes matching (IAM). Next, we study the impacts of the iterative attributes matching mechanism by removing it from Editable-DeepSC. As shown in Figure 10, when the iterative attributes matching mechanism is excluded, Editable-DeepSC will be unable to precisely perform the anticipated editings according to the given texts in a fine-grained way. This again provides evidence for the contribution of our IAM mechanism.
The influences of model fine-tuning. Similarly, we also investigate the impacts of the model fine-tuning mechanism in Table 4, Table 5, and Figure 11. We discover that under the extreme channel condition of 6-6 dB SNR, the fine-tuning mechanism plays a crucial role in preserving the quality and fidelity of edited images, while removing it will conversely generate unnatural or even collapsed regions. These results also verify the contribution of our fine-tuning mechanism. Moreover, our adoption of $r_{ft}=6$ yields a satisfactory trade-off between fine-tuning effects and GPU energy costs.
Table 6: GPU energy costs ($\downarrow$) of different methods averaged over each image under different resolutions.
@@ 5.5 Further Analysis
Table 8: Quantitative results when adopting different variants of the channel decoder (CD) on the CelebA dataset at the SNR of 6-6 dB.
Model Variant LPIPS ($\downarrow$) FID ($\downarrow$) KID ($\downarrow$) CLIP ($\uparrow$) Editable-DeepSC 0.1616 66.64 -0.0012 0.2439 w/ weak CD 0.2488 91.54 0.0102 0.2315 w/o CD 0.2965 103.81 0.0330 0.2305
Table 9: Estimated mutual information (MI) between the original distribution $K_{0}$ and the intermediate distributions at different processing stages. The decreasing MI along the processing chain aligns well with the Data Processing Inequality (DPI).
Distribution Pair $K_{0},K_{1}$ $K_{0},K_{2}$ $K_{0},K_{3}$ $K_{0},K_{4}$ Mutual Information 2.6972 2.4683 1.8304 1.6627
Table 10: Component contribution analysis under varying channel conditions. We report the CLIPsim ($\uparrow$) values of different model variants on the CelebA dataset.
Model Variant 6-6 dB 0 dB $6$ dB $12$ dB $18$ dB Editable-DeepSC 0.2439 0.2454 0.2458 0.2476 0.2487 w/ weak GAN inversion 0.2296 0.2329 0.2357 0.2371 0.2388 w/o IAM 0.2266 0.2304 0.2312 0.2297 0.2310 w/o FT 0.2213 0.2242 0.2318 0.2362 0.2366
Table 11: Inference speed comparison on the CelebA dataset. The time costs of GAN inversion and JECC iterations are also provided.
Method Average Time (s) DeepSC+Talk2Edit 35.16 JPEG+LDPC+Talk2Edit 36.36 Talk2Edit+DeepSC 35.26 Talk2Edit+JPEG+LDPC 35.68 GAN Inversion (Ours) 7.79 JECC Iterations (Ours) 0.35 Editable-DeepSC 10.78
@@
Ablation study on the channel decoder (CD) mechanism. In Table 8, we compare three variants: the full Editable-DeepSC model, a version where the channel decoder is pre-trained at a high SNR but tested at 6-6 dB (i.e., w/ weak CD), and a version where the channel decoder is completely removed (i.e., w/o CD). The results show that removing the channel decoder leads to the worst performance across all metrics, as the model performs semantic decoding directly on the severely corrupted received signal $Y$ without any error correction. The weak CD variant shows intermediate performance, indicating that a channel decoder trained under a mismatched SNR condition still provides partial robustness, but is still significantly less effective than the properly fine-tuned decoder in the full model. These findings demonstrate the critical role of the SNR-aware channel decoder in mitigating semantic errors induced by channel noise.
Quantitative computation of mutual information (MI). In Section 3.3, we emphasize that the separate schemes introduce more data processing steps, which leads to the loss of semantic mutual information. To quantitatively validate this theoretical analysis, we employ the Mutual Information Neural Estimation (MINE) method [3] to estimate the MI between the original image distribution $K_{0}$ and the intermediate distribution $K_{s}$ at different processing stages. The results are summarized in Table 9. We find that the MI values exhibit a monotonic decrease along the processing chain: $\mathcal{I}(K_{0};K_{1})>\mathcal{I}(K_{0};K_{2})>\mathcal{I}(K_{0};K_{3})>\mathcal{I}(K_{0};K_{4})$. This trend aligns well with the Data Processing Inequality (DPI). By integrating editing into the communication chain and reducing the number of separate steps, our proposed JECC framework effectively preserves more semantic mutual information about the original task, which is the fundamental reason for its superior performance over the separate baselines.
Component contribution analysis under varying channel conditions. Table 10 quantitatively analyzes the contribution of each key component under varying channel conditions. The full Editable-DeepSC model consistently achieves the highest CLIPsim across all SNR levels, demonstrating the synergistic effect of its integrated design. The variant with weakened GAN inversion (adopting $r_{inv}=70$) shows a consistent performance drop, indicating that high-fidelity semantic compression is foundational for preserving editing semantics. The most significant performance degradation, especially at medium-to-high SNRs, occurs when removing the Iterative Attributes Matching (IAM) mechanism, highlighting its critical role in achieving precise semantic alignment with text instructions. Conversely, the variant without fine-tuning (FT) suffers the most severe performance loss at low SNRs (e.g., 6-6 dB), underscoring that the SNR-aware adaptation module is primarily responsible for maintaining robustness in harsh channel conditions. In summary, GAN inversion provides the essential compressed representation, the JECC mechanism enables accurate editing, and fine-tuning ensures channel resilience. Their combination in Editable-DeepSC optimally balances these requirements for reliable cross-modal semantic communications.
Inference speed analysis. From Table 11, we find that the complete Editable-DeepSC framework is significantly faster than all separate baselines. The time of Editable-DeepSC is primarily consumed by the GAN inversion stage (7.79 seconds), which is an optimization-based process to find the optimal latent representation. The JECC iterative editing stage is highly efficient (only 0.35 seconds on average). While the current implementation does not achieve millisecond-level latency for strict real-time video streaming, the total processing time of around 10 seconds is highly suitable for interactive applications such as social media photo editing or cloud-based portrait enhancement, where users expect to wait several seconds for high-quality, personalized results. This demonstrates a favorable trade-off between the exceptional bandwidth efficiency (as shown in Table 2 and Table 3) and the achievable processing speed for interactive scenarios.
@@ 6 Conclusion
In this paper, we propose Editable-DeepSC, a novel cross-modal semantic communication approach that tackles the communication challenges under dynamic facial editing scenarios. Specifically, we first theoretically analyze different transmission schemes that separately handle communications and editings. We find that they actually increase the data processing procedures and are more likely to lose semantic mutual information. In light of this, we emphasize the necessity to integrate editings into the communication chain, jointly optimize these two aspects, and preserve more semantic mutual information. To compress the high-dimensional data, we leverage inversion methods based on pre-trained StyleGAN priors for semantic coding. To tackle the dynamic channel noise conditions, we propose SNR-aware channel coding based on model fine-tuning, which introduces two lightweight trainable adapters to capture the distribution of new noise conditions. Extensive experiments demonstrate that compared to existing methods, Editable-DeepSC exhibits superior editing effects while significantly saving the bandwidth under multiple settings.
Limitations and Feature Work. While Editable-DeepSC demonstrates significant advantages in various settings, we acknowledge the practical deployment challenges for resource-constrained mobile transmitters. The iterative closed-loop checks (e.g., StyleGAN generation and attribute prediction) indeed impose computational burdens. For future work, we hope to explore deployment strategies such as model distillation for lightweight edge deployment [16] or edge-cloud collaborative computing [49], enhancing practicality while preserving semantic communication benefits.
Acknowledgements.
This work is supported in part by the National Natural Science Foundation of China under grant 62571298, 62576122, 62301189, and Shenzhen Science and Technology Program under Grant KJZD20240903103702004.